OpenStudy (anonymous):

Which is an equation of the circle with a radius of 9 units and center at (–4, 2)?

OpenStudy (anonymous):

Circle equation is (x-a)^2 + (y-b)^2 = r^2 (a,b) will be the center and r the radius.

OpenStudy (anonymous):

its one of these... A. (x − 9)2 + (y + 4)2 = 4 B. (x − 4)2 + (y + 2)2 = 81 C. (x + 4)2 + (y − 2)2 = 81 D. (x − 2)2 + (y + 4)2 = 81

OpenStudy (anonymous):

C

OpenStudy (anonymous):

u rock